# Statistical Taylor Expansion: A New and Path-Independent Method for Uncertainty Analysis 111

### v11

@@ 2.11 Dependency Tracing
Variance arithmetic employs dependency tracing to ensure that the calculated mean and variance fit statistics rigorously. However, dependency tracing comes with a price: variance calculations are generally more complex than value calculations, exhibiting narrower convergence range for input variables. Dependency tracing also implies that the results of statistical Taylor expansion should remain path independent.
@@ 2.12 Traditional Execution and Dependency Problem
Dependency tracing requires an analytic form of the function to apply statistical Taylor expansion for the result mean and variance, as in Formula (2.6), (2.3), (2.9), and (2.3). This requirement often conflicts with conventional numerical methods for on analytic functions:
  • Traditionally, intermediate variables are widely used in computations; however, this practice disrupts dependency tracing by obscuring the relationships between original input variables.
  • Similarly, conditional executions are commonly employed to optimize performance and minimize rounding errors, such as using Gaussian elimination to minimize floating-point rounding errors in matrix inversion [37]. For dependency tracing, such conditional executions should be replaced by direct matrix inversion as described in Section 6.
  • Furthermore, traditional approaches frequently apply approximations to the result values during executions. Under the statistical Taylor expansion, Formula (2.3) shows that the variance converges more slowly than the value in Taylor expansion. Consequently, approximation strategies should prioritize variances than values. Section 6 illustrates this with a first-order approximation for computing a matrix determinant.
  • Traditionally, results from mathematical library functions are accepted without question, with accuracy assumed down to the last bit. As demonstrated in Section 8, statistical Taylor expansion enables the detection of numerical errors within these functions and necessitates recalculating them with uncertainty explicitly included in the output.
  • In conventional practice, an analytic expression is often decomposed into simpler, ostensibly and independent arithmetic operations such as negation, addition, multiplication, and division  [^5] Sometimes square root is also regarded as an arithmetic operation.. However, this decomposition introduces dependency problems in floating-point arithmetic, interval arithmetic, and statistical Taylor expansion. For example, if $x^{2}-x$ is calculated as $x^{2}-x$, x(x1)x(x-1), and $(x-\frac{1}{2})^{2}-\frac{1}{4}$, only $(x-\frac{1}{2})^{2}-\frac{1}{4}$ gives the correct result, while the other two give wrong results for wrong independence assumptions between $x^{2}$ and $x$, or between x1x-1 and $x$, respectively.
@@
Dependency tracing thus removes nearly all flexibility in traditional numerical executions, eliminating associated dependency problems. Consequently, all conventional numerical algorithms require reevaluation or redesign to align with the principles of statistical Taylor expansion.
@@ 3 Variance Arithmetic
This paper confines the calculating of variances to ideal variances $\widehat{\delta^{2}f}$ when $\kappa=5$, and assumes that that the input distribution is Gaussian, although variance arithmetic also implements statistical Taylor expansion for other cases. Furthermore, for discussion simplicity, Taylor coefficients in Formula (2.5) and (2.8) are all assumed to be precise.
@@ 3.1 Numerical Representation
Variance arithmetic represents an imprecise value $x\pm\delta x$ using a pair of 64-bit standard floating-point numbers. Other conventional numerical numbers need to convert to this format.
If the least $20$ significant bits of its significand are all 0 in a standard floating-point number, the value is considered precise, representing a 2’s fractional with a possibility no less than $1-2^{-20}=1-2.384\;10^{-7}$. Otherwise, a standard floating-point value is considered imprecise with the uncertainty given as $1/\sqrt{3}$-fold of the ULP of the value, where ULP stands for Unit in the Last Place for a conventional floating-point number [9]. This follows from the fact that the pure rounding error of round-to-nearest is uniformly distributed within half bit of the significand of the floating-point value [1].
If an integer number is within the range $2^{53}+1,+2^{53}-1]$ of the significand of a 64-bit standard floating-point number, its uncertainty is zero. Otherwise, it is first converted to a conventional floating-point value before being converted to an imprecise value.
3.2 Calculation Rules
Variance arithmetic implements statistical Taylor expansion with several practical assumptions. Due to finite precision and range of conventional floating-point representation, $\zeta(n)$ can be computed only for $n<448$ before becoming infinite under the IEEE 64-bit floating-point format, Consequently, the following numerical rules are introduced:
  • finite: The result must remain finite.
  • monotonic: As a necessary condition for convergence, the last $20$ terms of the expansion must decrease monotonically in absolute value, ensuring that the probability of the expansion to an absolute increase is no more than $2^{-20}=9.53\;10^{-7}$. Unless all the remaining terms in the expansion are known to be precise zeros, an expansion is always executed to the full 448 terms for the monotonic check.
  • positive: At any order, the expansion variance must be positive.
  • stable: For sufficiently fast convergence, the value of the last expansion term must be less than the $\epsilon(5)=5.73\;10^{-7}$ times of both the result uncertainty and the result absolute value.
  • reliable: At any order, the value of the variance must be less than $1/\kappa$-fold of the uncertainty of the variance.
The validity of these numerical rules of variance arithmetic will be demonstrated in subsequent sections.
3.2.1 Finite
@@ 3.2.2 Monotonic
From Formula (2.8), the convergence condition applied to $\log(x\pm\delta x)$ is $P(x)\leq 1/\kappa=1/5$, which is numerically confirmed as $P(x)\lesssim 0.20086$. Beyond this upper bound, the expansion is no longer monotonic. Variance arithmetic rejects the distributional zero of $\log(x)$ in the range of $x-\delta x,x+\delta x]$ statistically due to the divergence of Formula (2.29) mathematically, with $\zeta(2n)$ as the connection of these two perspectives.
@@
From Formula (2.44), and except when $c$ is a natural number, Formula (2.35) for $(x\pm\delta x)^{c}$ converges near $P(x)\simeq 1/\kappa=1/5$, with the upper bound $P(x)$ increasing with $c$. This trend is approximately confirmed in Figure 7, and beyond the upper bound $P(x)$, the expansion is no longer monotonic. Qualitatively, $\delta^{2}1/x$ converges more slowly than $\delta^{2}\sqrt{x}$, consistent with Formula (2.42) and (2.41).
When the input uncertainty is Uniformly distributed, Figure 8 illustrates that the input uncertainty upper bound is close to $1/\sqrt{3}$ according to Formula (2.44). The bounded momentum $\zeta(2n)$ for Uniform distribution is much less than that for Normal distribution, such the expansion in Formula (2.44) contains $652$ terms instead of $448$ terms. The upper bounds of input uncertainty in Figure 7 and 8 looks identical, however, the resulting uncertainty biases and uncertainties in Figure 7 are $3$ order-of-magnitude smaller than those in Figure 8 when the exponent is less than 0. This suggests that convergence is not the only consideration in statistical Taylor expansion because uncertainty bias or uncertainty can also invalidate a result in practice.
3.2.3 Positive
@@
In some cases, the variance expansion may produce negative results, as in Formula (2.32) for $sin(x\pm\delta x)$. Figure 9 demonstrates that the upper bound of $\delta x$ for $sin(x\pm\delta x)$ varies periodically between $0.318\pi$ and $0.416\pi$, which is less than $\pi$ of $\sin(x)$. Beyond this upper bound, the expansion is no longer positive in Figure 9.
@@ 3.2.5 Reliable
3.3 Continuity
In variance arithmetic, the result mean, variance and histogram are generally continuous across parameter space. For example, $\delta x$ has an upper bound for $(x\pm\delta x)^{c}$ to converge except when $c$ is a natural number. The result mean, variance and histogram of $(x\pm\delta x)^{c}$ remains continuous around $c=n$. Table 1 shows that the result of $(1\pm\delta x)^{n\pm d},0<d\ll 1$ is very close to that of $(1\pm\delta x)^{n}$, even though the former had upper bound for $\delta x$, while the latter does not.
A statistical bounding range in variance arithmetic can include a distributional pole if the analytic function is defined in its vicinity. The presence of such poles does not disrupt continuity in of the result mean, variance, or histogram. Figure 12 illustrates the histograms of $(x\pm 0.2)^{n}$ when x=0,0.2,+0.2x=0,-0.2,+0.2 and $n=2,3$.
  • In each case, the transition from $x=0$ to $x=0.2$ is continuous.
A statistical bounding range inf variance arithmetic cannot encompass more than one distributional pole, as this condition causes the corresponding statistical Taylor expansion becomes negative. An illustrative example is $\sin(x)$ in Figure 9.
$n\pm d$ $0\pm 10^{-6}$ $1\pm 10^{-6}$ $2\pm 10^{-6}$ $3\pm 10^{-6}$
Upper Bound $\delta x$ $0.2006$ $0.2014$ $0.2018$ $0.2020$
3.4 Tracking Rounding Error
Figure 12: Residual error of $\sum_{j=0}^{224}x^{j}-\frac{1}{1-x}$ vs $x$ (x axis). The y-axis to the left shows both the value and the uncertainty of the residual errors. The y-axis to the right indicates the expansion order needed to reach stable value for each $x$.
$\sum_{j=0}^{N}c_{j}(x+\tilde{x})^{j}=\sum_{j=0}^{N}\tilde{x}^{j}\sum_{k=0}^{N-j}x^{k-j}c_{j+k}\begin{pmatrix}j+k\\ j\end{pmatrix}$; (3.1)
Figure 12 illustrates the residual error of $\sum_{j=0}^{224}x^{j}-\frac{1}{1-x}$, where the polynomial $\sum_{j=0}^{224}x^{j}$ is computed using Formula (3.1), $\frac{1}{1-x}$ is computed using Formula (2.33), and $x$ is initiated as a floating-point value. Because $\eta(2n)$ is limited to $2n\leq 448$, Formula (3.1) for polynomial evaluation is restricted to $N\leq 224$, so that $\sum_{j=0}^{224}x^{j}$ has lower expansion order than that of $\frac{1}{1-x}$. Figure 12 shows:
  • When $x\in[-0.73,0.75]$, the required expansion order is no more than $224$, meaning that the residual error reflects solely the rounding error between $\sum_{j=0}^{224}x^{j}$ and $\frac{1}{1-x}$. A detailed analysis indicates that the maximal residual error is $4$ times the ULP of $\frac{1}{1-x}$. The calculated uncertainty bounds the residual error effectively for all $x\in[-0.73,0.75]$.
  • When $x\not\in[-0.74,+0.75]$, for example when x=0.75,0.76x=-0.75,0.76, the required expansion order exceeds $224$, so that the residual error arises from the insufficient expansion order of $\sum_{j=0}^{224}x^{j}$. The residual error magnitude increases as $|x|\rightarrow 1$, reaching about $50$ when $x=0.98$.
@@ 3.5 Comparison
Two imprecise values can be compared statistically for their difference.
When the value difference is zero, the two imprecise values are equal. In statistics, such two imprecise values have $50\%$ possibility to be less than or greater to each other but no chance to be equal to each other [5]. In variance arithmetic, however, they are treated as neither less nor greater than each other, thus they are equal.
Otherwise, the standard z-statistic method [5] is applied to determine whether they are equal, less than, or more than each other. For example, the difference between $1.002\pm 0.001$ and $1.000\pm 0.002$ is $0.002\pm 0.00224$, yielding $z=0.002/0.00224$, and the probability for them to be not equal is $\xi(|z|/\sqrt{2})=62.8\%$, in which $\xi(z)$ is the cumulative density function for Normal distribution [5]. If the threshold probability for inequality is $50\%$, then $1.000\pm 0.002<1.002\pm 0.001$. Alternatively, an equivalent bounding range for z can be used, such as $|z|\leq 0.67448975$ for an equal probability threshold of $50\%$.
4 Verification of Statistical Taylor Expansion
@@ 4.1 Verification Methods and Standards
Statistical Taylor expansion is verified through the validation of variance arithmetic. Analytic functions or algorithms with precisely known results are used to compare the outputs of variance arithmetic using the following statistical properties:
@@
One objective of uncertainty-based calculation is to account precisely for all input errors from every source, thereby achieving ideal coverage:
@@
  3. 3.
    The uncertainty response should match the expected functional form. For a linear function, the output uncertainties should increase linearly with input uncertainties, with the ratio of the output to input uncertainty means defined as uncertainty response ratio.
  4. 4.
    The calculation response should follow the expected trend; for instance, increasing the amount of calculations should result in larger uncertainties.
If the precise result is unknown, as outlined in Section 6, the result error distribution can be used to determine whether ideal coverage is achieved.
If however, the input uncertainty is only accurate to the order of magnitude, the proper coverage is obtained when the error deviations fall within the range of $[0.1,10]$.
When an input contains unspecified errors, such as numerical errors in library functions, Gaussian noise with progressively increasing deviations can be added, until ideal coverage is attained. The minimal noise deviation required provides a good estimation of the magnitude of unspecified input uncertainties. The presence of ideal coverage is a necessary verification step to ensure that Formula (2.3) or (2.3) have been applied correctly in the given context. The input noise range yielding ideal coverage defines the ideal application range for input uncertainties.
Gaussian noises of varying deviations can also be added to the input to evaluate a function’s uncertainty response.
@@ 4.2 Types of Uncertainties
  • Input uncertainties: The examples in this paper demonstrate that when input uncertainty precision is $10^{-15}$ or larger, variance arithmetic can achieve ideal coverage for input uncertainties.
  • Rounding errors: Empirical results indicate that variance arithmetic can provide proper coverage for rounding errors.
  • Truncation errors: Variance arithmetic avoids truncation errors with its stable rule.
  • External errors: External errors are the value errors which are not specified in input uncertainties, such as numerical errors in library functions. Section 8 examines the effects of numerical errors in $\sin$ and $\tan$, showing that when these external errors are sufficiently large enough, neither ideal coverage nor proper coverage can be achieved. This finding indicates that library functions must be recalculated to include the corresponding uncertainty of each value.
  • Modeling errors: Modeling errors arise when an approximate analytic solution is used, or when a real-world problem is simplified to obtain a solution. For example, the predecessor of statistical Taylor expansion [1] demonstrated that the discrete Fourier transformation is only an approximation for the mathematically defined continuous Fourier transformation, containing modeling errors. Conceptually, modeling errors originate in mathematics and are therefore outside the domain of statistical Taylor expansion.
@@ 4.3 Types of Calculations to Verify
Algorithms of distinct nature each representative of its category are required to test the broad applicability of variance arithmetic [1]. An algorithm can be categorized by comparing the amount of its input and output data as [1]:
@@
A transformation algorithm produces output data of approximately the same quantity as its input, with the overall information contained remaining largely unchanged. For reversible transformations, a unique requirement is to recover the original input uncertainties after a round-trip transformation: performing a forward transformation followed by its reverse transformation. DFT (discrete Fourier transformation) is a typical reversible transformation algorithm: it has the same amount of input and output data, and its output can be transformed back to the input using essentially the same process. A test of variance arithmetic using FFT (fast Fourier transformation, which is an implementation of DFT) algorithms is presented in Section 8.
@@
A generation algorithm produces substantially more output data than input data. Such algorithms encode mathematical knowledge into data. Some generation algorithms are theoretical calculations which involve no imprecise input, so all resulting uncertainty arises from rounding errors. Section 9 presents a generation algorithm, which generates a sine function table using trigonometric relations and two precise inputs: $sin(0)=0$ and $sin(\pi/2)=1$.
A reduction algorithm yields significantly fewer output data than input data, as in numerical integration, or statistical characterization of a data set. In this process, certain information is lost while other information is extracted. Conventional wisdom is that a reduction algorithm generally benefits from a larger input data set [5]. A test of statistical characterization through sampling is described in Section 5.

### v12

@@ 2.11 Dependency Tracing
Variance arithmetic employs dependency tracing to ensure that the calculated mean and variance satisfy statistics rigorously. However, dependency tracing comes at a cost: variance calculations are generally more complex than value calculations and exhibits a narrower convergence range for input variables. Dependency tracing also implies that the results of statistical Taylor expansion must remain path independent.
@@ 2.12 Traditional Execution and Dependency Problem
Dependency tracing requires an analytic form of the function to apply statistical Taylor expansion for the result mean and variance, as in Formula (2.9), (2.3), (2.12), and (2.3). This requirement often conflicts with conventional numerical methods for analytic functions:
  • Traditionally, intermediate variables are widely used in computations; however, this practice disrupts dependency tracing by obscuring the relationships among the original input variables.
  • Similarly, conditional executions are often employed to optimize performance and minimize rounding errors, for example, using Gaussian elimination to minimize floating-point rounding errors in matrix inversion [37]. For dependency tracing, such conditional executions should instead be replaced by direct matrix inversion as described in Section 7.
  • Furthermore, traditional approaches frequently apply approximations to result values during execution. Under the statistical Taylor expansion, Formula (2.3) shows that the variance converges more slowly than value in statistical Taylor expansion. Consequently, approximation strategies should prioritize variances than values. Section 7 illustrates this principle through a first-order approximation used in computing a matrix determinant.
  • Traditionally, results from mathematical library functions are accepted without scrutiny, with accuracy assumed down to the last bit. As demonstrated in Section 9, statistical Taylor expansion enables the detection of numerical errors within these functions and requires that they be recalculated with uncertainty explicitly incorporated into the output.
  • In conventional practice, an analytic expression is often decomposed into simpler, ostensibly and independent arithmetic operations such as negation, addition, multiplication, division, square root, and library calls. However, this decomposition introduces dependency problems in floating-point arithmetic, interval arithmetic, and statistical Taylor expansion. For example, if $x^{2}-x$ is calculated as $x^{2}-x$, x(x1)x(x-1), and $(x-\frac{1}{2})^{2}-\frac{1}{4}$, only $(x-\frac{1}{2})^{2}-\frac{1}{4}$ gives the correct result, while the other two give wrong results for wrong independence assumptions between $x^{2}$ and $x$, or between x1x-1 and $x$, respectively.
@@
Dependency tracing therefore removes nearly all flexibility from traditional numerical executions, effectively eliminating the associated dependency problems. Consequently, all conventional numerical algorithms must be reevaluated or redesigned to align with the principles of statistical Taylor expansion.
@@ 3 Variance Arithmetic
Variance arithmetic implements statistical Taylor expansion. Because of the finite precision and limited range of conventional floating-point representation, $\zeta(n)$ can only be computed to limited terms. Consequently, the following numerical rules are introduced:
  • finite: The resulting value and variance must remain finite.
  • monotonic: As a necessary condition for convergence, the last $20$ terms of the expansion must decrease monotonically in absolute value, ensuring that the probability of the expansion exhibiting an absolute increase is no more than $2^{-20}=9.53\;10^{-7}$. Unless all the remaining terms in the expansion are known to be precisely zeros, each expansion is executed to the full 448 terms for the monotonicity check.
  • positive: At every order, the expansion variance must be positive.
  • stable: To avoid truncation error [13], the value of the last expansion term must be less than $5.73\;10^{-7}$ times of both the result uncertainty and the result absolute value. This rule ensures sufficiently fast convergence in the context of monotonic convergence.
  • reliable: At every order, the uncertainty of the variance must be less than $1/5$ times of the value of the variance.
For simplicity of discussion, This paper confines the calculation of variances to ideal variances $\widehat{\delta^{2}}f$ and assumes that the input distribution is Gaussian with $\kappa_{s}=5$. Furthermore,the Taylor coefficients in Formula (2.8) and (2.11) are assumed to be precise.
@@ 3.1 Numerical Representation
Variance arithmetic represents an imprecise value $x\pm\delta x$ using a pair of 64-bit standard floating-point numbers. All other conventional numerical numbers must be converted to this format.
If the least $20$ significant bits of the significand in a standard floating-point number are all 0, the value is considered precise, representing a 2’s fractional with a probability no less than $1-2^{-20}=1-2.384\;10^{-7}$. Otherwise, a standard floating-point value is considered imprecise with its uncertainty defined as $1/\sqrt{3}$ times of the ULP of the value, where ULP refers to the Unit in the Last Place in conventional floating-point representation [9]. This follows from the fact that the pure rounding error in round-to-nearest mode is uniformly distributed within half bit of the significand of a floating-point value [1].
If an integer number is within the range $2^{53}+1,+2^{53}-1]$ of the significand of a 64-bit standard floating-point number, its uncertainty is zero. Otherwise, it is first converted to a conventional floating-point value before being transformed into an imprecise value.
Variance arithmetic uses floating-point arithmetic for computation.
3.2 Finite
@@ 3.3 Monotonic
From Formula (2.8), the convergence condition applied to $\log(x\pm\delta x)$ is $P(x)\leq 1/\kappa=1/5$, which is numerically confirmed as $P(x)\lesssim 0.20086$. Beyond this upper bound, the expansion is no longer monotonic. Variance arithmetic rejects the distributional zero of $\log(x)$ in the range of $x-\delta x,x+\delta x]$ statistically due to the divergence of Formula (2.31) mathematically, with $\zeta(2n)$ providing the connection between these two perspectives.
@@
From Formula (2.46), and except when $c$ is a natural number, Formula (2.37) for $(x\pm\delta x)^{c}$ converges near $P(x)\simeq 1/\kappa=1/5$, with the upper bound $P(x)$ increasing with $c$. This trend is approximately confirmed in Figure 9, and the expansion is no longer monotonic beyond the upper bound $P(x)$. Qualitatively, $\delta^{2}1/x$ converges more slowly than $\delta^{2}\sqrt{x}$, consistent with Formula (2.44) and (2.43).
When the input uncertainty is Uniformly distributed, Figure 10 illustrates that the input uncertainty upper bound is close to $1/\kappa=/\sqrt{3}$ according to Formula (2.46). The bounded momentum $\zeta(2n)$ for Uniform distribution is much less than that for Normal distribution, such that the expansion contains $652$ terms instead of $448$ terms. The upper bounds of input uncertainty in Figure 9 and 10 looks identical, however, the resulting uncertainty biases and uncertainties in Figure 9 are $3$ order-of-magnitude smaller than those in Figure 10 when the exponent $c$ is less than 0. Even a result converges, it can still be abandoned due to too large uncertainty bias or uncertainty.
3.4 Positive
@@
In some cases, the variance expansion may yield negative results, as in Formula (2.34) for $sin(x\pm\delta x)$. Figure 11 shows that the upper bound of $\delta x$ for $sin(x\pm\delta x)$ varies periodically between $0.318\pi$ and $0.416\pi$. Beyond this upper bound, the expansion is no longer positive.
@@ 3.7 Comparison
Two imprecise values can be compared statistically based on their difference.
When the value difference is zero, the two imprecise values are considered equal. In statistics, such two values have a $50\%$ possibility of being either less than or greater to each other but zero probability of being exactly equal [5]. In variance arithmetic, however, they are treated as neither less nor greater than each other and therefore are considered equal.
Otherwise, the standard z-statistic method [5] is applied to determine whether two imprecise values are statistically equal, less than, or greater than each other. For example, the difference between $1.002\pm 0.001$ and $1.000\pm 0.002$ is $0.002\pm 0.00224$, yielding $z=0.002/0.00224$. The probability that they are not equal is $\xi(|z|/\sqrt{2})=62.8\%$, in which $\xi(z)$ is the cumulative distribution function for Normal distribution [5]. If the threshold probability for inequality is set at $50\%$, then $1.000\pm 0.002<1.002\pm 0.001$. Alternatively, an equivalent bounding range for z can be used, such as $|z|\leq 0.67448975$ for an equal probability threshold of $50\%$.
Because the result of comparison depends on threshold probability which is application specific, comparison is not part of variance arithmetic.
4 Verification of Variance Arithmetic
@@ 4.1 Verification Methods and Standards
Analytic functions or algorithms with precisely known results are used to evaluate the outputs of variance arithmetic based on the following statistical properties:
@@
One objective of uncertainty-based computation is to account precisely for all input errors from every source, thereby achieving ideal coverage:
@@
If the precise result is unknown, the resulting normalized error distribution can be used to assess whether ideal coverage is achieved.
However, if the input uncertainty is known only to order of magnitude, proper coverage is achieved when the error deviations fall within the range $[0.1,10]$.
When an input contains unspecified errors, such as numerical errors in library functions, Gaussian noise with progressively increasing deviations can be added, until ideal coverage is attained. The minimal noise deviation required provides a good estimate of the magnitude of the unspecified input uncertainties. Achieving ideal coverage serves as a necessary verification step to ensure that Formula (2.3) or (2.3) have been applied correctly within the given context. The input noise range that yields ideal coverage defines the ideal application range for input uncertainties.
Besides coverage, other validation criteria are also used:
  • The uncertainty response should match the expected functional form. For a linear function, the output uncertainties should increase linearly with the input uncertainties, with the ratio of output to input uncertainty means defined as uncertainty response ratio.
  • The calculation response should follow the expected trend; for instance, increasing the number of calculations should result in larger output uncertainties.
Noises of varying deviations can also be added to the input to evaluate a function’s uncertainty response.
@@ 4.2 Types of Uncertainties
  • Input uncertainties: The examples presented in this paper demonstrate that when the precision of input uncertainties is $10^{-15}$ or larger, variance arithmetic can achieve ideal coverage for input uncertainties.
  • Rounding errors: Empirical results indicate that variance arithmetic provides proper coverage for rounding errors.
  • Truncation errors: Variance arithmetic avoids truncation errors with its stable rule. However, using Formula (5.1) for polynomial expansion can result in truncation errors when the choice of the expansion order is insufficient, as illustrated in Figure 14 .
  • External errors: External errors are value errors not specified in the input uncertainties, such as numerical errors in library functions. Section 9 examines the effects of numerical errors of library sine function, showing that when these external errors are sufficiently large, neither ideal coverage nor proper coverage can be achieved. This finding indicates that library functions must be recalculated to explicitly include the corresponding uncertainty for each computed value.
  • Modeling errors: Modeling errors arise when an approximate analytic solution is used, or when a real-world problem is simplified to make a solution tractable. For example, Section 9 demonstrates that the discrete Fourier transform (DFT) is only an approximation of the mathematically defined continuous Fourier transform (FT), and therefore contains modeling errors. Conceptually, modeling errors originate in mathematics within and are thus outside the domain of statistical Taylor expansion.
@@ 4.3 Types of Calculations to Verify
Algorithms of distinct natures, with each representative of its respective category, are required to test the broad applicability of variance arithmetic [1]. An algorithm can be categorized by comparing the amount of its input and output data as [1]:
@@
A transformation algorithm produces output data of approximately the same quantity as its input, with the overall information content remaining largely unchanged. For reversible transformations, a unique requirement is to recover every original input for both value and uncertainty after a round-trip transformation that is performing a forward transformation followed by its reverse transformation. The discrete Fourier transform (DFT) is a typical reversible transformation algorithm: it has the same amount of input and output data, and its output can be transformed back into the input using essentially the same process. A test of variance arithmetic using the fast Fourier transform (FFT, which is an implementation of DFT) algorithms is presented in Section 9.
@@
A generation algorithm produces substantially more output data than input data. Such algorithms encode mathematical knowledge into data. Certain generation algorithms are purely theoretical calculations that involve no imprecise input, so all resulting uncertainty arises solely from rounding errors. Section 10 presents a generation algorithm that generates a sine function table using trigonometric relations and two precise inputs: $sin(0)=0$ and $sin(\pi/2)=1$.
A reduction algorithm yields significantly fewer output data than input data, as in numerical integration, or in the statistical characterization of a data set. In this process, certain information is lost while other information is extracted. As a statistical approach, variance arithmetic inherently carries out statistical reduction. For example, the result of averaging $N$ imprecise values with each uncertainty close to $\delta x$ is $\overline{x}\pm\frac{\delta x}{\sqrt{N}}$, reflecting central limit theorem [5] naturally.
5 Polynomial
Formula (5.1) presents polynomial Taylor expansion:
$\sum_{j=0}^{N}c_{j}(x+\tilde{x})^{j}$ $=\sum_{j=0}^{N}\tilde{x}^{j}P_{j},\;\;\;P_{j}\equiv\sum_{k=0}^{N-j}x^{k-j}c_{j+k}\begin{pmatrix}j+k\\ j\end{pmatrix}$; (5.1)
5.1 Tracking Rounding Error
Figure 13: Residual error of $\sum_{j=0}^{224}x^{j}-\frac{1}{1-x}$ vs $x$ (x-axis). The y-axis to the left shows both the value and the uncertainty of the residual errors. The y-axis to the right indicates the expansion order needed to reach stable value for each $x$.
Figure 14 shows the residual error of $\sum_{j=0}^{224}x^{j}-\frac{1}{1-x}$, where the polynomial $\sum_{j=0}^{224}x^{j}$ is computed using Formula (5.1), $\frac{1}{1-x}$ is computed using Formula (2.35), and $x$ is initiated as a floating-point value. Because $\eta(2n)$ is limited to $2n\leq 448$, Formula (5.1) for polynomial evaluation is restricted to $N\leq 224$, so that $\sum_{j=0}^{224}x^{j}$ has lower expansion order than that of the statistical Taylor expansion of $\frac{1}{1-x}$. Figure 14 shows:
  • When $x\in[-0.73,0.75]$, the required expansion order is no more than $224$, which indicates that the residual error reflects solely the rounding error between $\sum_{j=0}^{224}x^{j}$ and $\frac{1}{1-x}$. A detailed analysis indicates that the maximal residual error is $4$ times the ULP of $\frac{1}{1-x}$. The calculated uncertainty bounds the residual error effectively for all $x\in[-0.73,0.75]$.
  • When $x\not\in[-0.74,+0.75]$, the required expansion order exceeds $224$, so that the residual error arises from the insufficient expansion order of $\sum_{j=0}^{224}x^{j}$. The residual error magnitude increases as $|x|\rightarrow 1$, reaching approximately $50$ when $x=0.98$.
5.2 Continuity
$n\pm d$ $0\pm 10^{-6}$ $1\pm 10^{-6}$ $2\pm 10^{-6}$ $3\pm 10^{-6}$
Upper Bound $\delta x$ $0.2006$ $0.2014$ $0.2018$ $0.2020$
In variance arithmetic, the result mean, variance and histogram are generally continuous across parameter space. For example, $\delta x$ has an upper bound for $(x\pm\delta x)^{c}$ to converge except when $c$ is a natural number. The result mean, variance and histogram of $(x\pm\delta x)^{c}$ remain continuous around $c=n$. Table 2 shows that the result of $(1\pm\delta x)^{n\pm d}$ where $0<d\ll 1$ is very close to that of $(1\pm\delta x)^{n}$, even though the former has an upper bound for $\delta x$, while the latter does not.
A statistical bounding range in variance arithmetic can include a distributional pole if the analytic function is defined in its vicinity. The presence of such poles does not disrupt the continuity of the result mean, variance, or histogram. Figure 14 illustrates the histograms of $(x\pm 0.2)^{n}$ when x=0,0.2,+0.2x=0,-0.2,+0.2 and $n=2,3$.
In each case, the transition from $x=0$ to $x=0.2$ is continuous.
A statistical bounding range in variance arithmetic cannot encompass more than one distributional pole, as this condition causes the corresponding statistical Taylor expansion becomes negative. An illustrative example is $\sin(x)$ shown in Figure 11.